OpenStudy (anonymous):

If f(x) = xsquared does f(a+1) = (a+1)squared

OpenStudy (anonymous):

what's the question?

OpenStudy (anonymous):

if f(x) = X squared does it follow that f(a+1) = (a+1) squared

OpenStudy (anonymous):

True :D

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

explain why

OpenStudy (anonymous):

replace x with (a+1).

OpenStudy (anonymous):

How do you evaluate the expresssion f(g(4)) when f(x)= 3x and g(x)= x-2

OpenStudy (anonymous):

f(g(4))=3(4-2)=6

OpenStudy (hba):

@firebird Do you understand ?

OpenStudy (anonymous):

no

OpenStudy (hba):

So let me explain that to you We have to find f(g(4) So we have to find g(4) first g(x)=x-2 then g(4)=

OpenStudy (anonymous):

note: if f(x)=3x, g(x)=x-2, then f(g(x))=3(g(x))=3(x-2)

OpenStudy (anonymous):

f(g(x)) :substitute g(x) in x of f(x).

OpenStudy (anonymous):

simply sub anything inside x as long it's in the domain.

OpenStudy (hba):

@firebird Do you follow me ?

OpenStudy (anonymous):

yes

OpenStudy (hba):

So what did you get ? g(4)=

OpenStudy (anonymous):

4-2=g(x)

OpenStudy (anonymous):

So under the same values of f(x) and g(x), g(f(4)) is equal to g(3(x)) which is equal to x-2(3x)?